### v3

# Tuning the Frequencies: Robust Training for Sinusoidal Neural Networks

Abstract
Sinusoidal neural networks have been shown effective as implicit neural representations (INRs) of low-dimensional signals, due to their smoothness and high representation capacity. However, initializing and training them remain empirical tasks which lack on deeper understanding to guide the learning process. To fill this gap, our work introduces a theoretical framework that explains the capacity property of sinusoidal networks and offers robust control mechanisms for initialization and training. Our analysis is based on a novel amplitude-phase expansion of the sinusoidal multilayer perceptron, showing how its layer compositions produce a large number of new frequencies expressed as integer combinations of the input frequencies. This relationship can be directly used to initialize the input neurons, as a form of spectral sampling, and to bound the network’s spectrum while training. Our method, referred to as TUNER (TUNing sinusoidal nEtwoRks), greatly improves the stability and convergence of sinusoidal INR training, leading to detailed reconstructions, while preventing overfitting.
1 Introduction
Sinusoidal multilayer perceptrons (MLPs) emerged as powerful implicit neural representations (INRs) for low-dimensional signals [26, 3, 35, 11]. In this context, the INR $f$ should fit the input data $\{x_{i},\mathscr{f}_{i}\}$ as close as possible, i.e. $f(x_{i})\approx\mathscr{f}_{i}$, without overfitting, thus encoding the signal implicitly in the MLP parameters. Therefore, two major properties are required: (1) $f$ needs high representation capacity to fit $\{x_{i},\mathscr{f}_{i}\}$; (2) $f$ should have bandlimit control to avoid frequencies bypassing the sampling rate.
Figure 1: We present TUNER, a robust and theoretically grounded training technique for sinusoidal MLPs, overcoming challenges in initialization and enabling bandlimiting control. Our experiments showcase TUNER’s strong initialization results against ReLU, FFM [28], and SIREN [26] (top), where all models use the same size and training conditions. TUNER achieves both fast and stable convergence (bottom-left) while reconstructing gradients without noise. We also compare with BACON [10] across two bandlimits (bottom-right), enhancing quality and avoiding ringing artifacts.
Training sinusoidal MLPs to satisfy the above properties is challenging, as their initialization and optimization process often lead to undesirable local minima [15]. Recent work made strides towards more effective learning of these models. For example, SIREN [26] proposed an initialization by projecting the input coordinates to a list of sines with frequencies randomly chosen in a range, similar to the Fourier feature mapping (FFM) approach [28]. This way, the model can reach high capacity, but may lead to overfitting with high frequencies resulting in noisy reconstructions. Defining an effective range for the bandlimit initialization remains mostly empirical and often results in noise, as the role of layer composition in generating frequencies is not fully understood. Additionally, uniform initialization may introduce undesired high frequencies and make it harder to model lower ones. More recently, BACON [10] proposes tighter bandlimit control by applying multiplicative filter networks (MFNs) [7] to limit the signal spectrum with a box filter, hard-truncating the spectrum. While effective in many cases, this may lead to ringing artifacts. Also, this technique lacks non-linear activations (MFNs are not neural networks), preventing the representation of fine details as efficiently as sinusoidal MLPs with similar sizes.
Overall, initializing and training sinusoidal MLPs remain empirical tasks which lack on deeper understanding to guide the learning. To approach these problems, we present TUNER (TUNing sinusoidal nEtwoRks), a training scheme for sinusoidal MLPs consisting of a robust initialization and bandlimit mechanism that greatly improves the capacity and convergence of INRs. In contrast to previous work, TUNER is grounded on a theoretical framework that guides our learning approach. We develop a new trigonometric identity (Thrm 1), resembling a Fourier series, and prove that such expansion produces a large number of frequencies in terms of the first layer (input frequencies). This explains the frequency generation process when composing layers, highlighting the importance of input layer initialization for improving capacity. Remarkably, the amplitudes of these frequencies are given by complex functions of the hidden weights, introducing a challenging training. To address this, we prove that those are bounded by a term depending only on the hidden weights (Thrm 2). We apply this result to control the bandlimit of the INR during training. Some of TUNER’s results are showcased in Fig. 1, in comparison to previous work, illustrating substantial enhancements to sinusoidal INR modeling. In summary, our contributions are:
  • We introduce a novel trigonometric identity that expands any hidden neuron into a wide sum of sines with frequencies being integer combinations of the input frequencies (Fig 2). The corresponding amplitudes, determined by the hidden weights, have a useful upper bound, offering effective tools for controlling the resulting signal.
  • The expansion enables a robust initialization scheme for sinusoidal INRs, resulting in a high capacity MLP that trains significantly faster than previous approaches. First, we initialize the input neurons, as they determine the INR spectrum. Next, we initialize the hidden weights, considering their role in setting the corresponding amplitudes.
  • We leverage the amplitude’s upper bound to control the bandlimit of a sinusoidal INR by designing schemes that bound the hidden weights during training. Together with our initialization, these bounds generate frequencies centered around the input, resulting in more stable training compared to previous approaches.
2 Related works
Implicit neural representations [30, 25, 8, 21, 32, 20] are a trending topic in machine learning, used to learn highly detailed signals in low-dimension domains. Current INR architectures use Fourier feature mappings [28] or sinusoidal activation functions [26] to bypass the spectral bias [17] common in ReLU multilayer perceptrons. The high representation capacity of sinusoidal INRs has motivated their use to represent a wide range of media objects. Examples include audio [6, 26], images [4, 27], face morphing [22], signed distance functions [23, 10, 26, 31, 13], displacement fields [33], surface animation [12, 14], multi-resolution signals [19, 10, 5, 29, 16], among others. Most of these exploit the derivatives of sinusoidal INRs in the loss functions.
Initialization. Considering sinusoidal activation functions in neural networks is a classical problem [18]; however, these INRs have been regarded as difficult to train [15]. Sitzmann et al. [26] overcome this by presenting a specific initialization scheme that allows training sinusoidal INRs, avoiding instability and ensuring convergence. Despite these advances, in practice, the initialization of such INRs remains an empirical task. In this work, inspired by Fourier series theory, we present a novel initialization method that allows us to train INRs with great convergence. Recently, several works have addressed the representation problem of sinusoidal INRs from different perspectives. Zell et al. [35] approached this problem by observing that the first layer of a sinusoidal INR is similar to a Fourier feature mapping. Here, in addition to improving the initialization of sinusoidal INRs, we present a training scheme for bounding the spectra of these networks.
Control of spectrum. One of the main drawbacks of sinusoidal INRs is the lack of frequency control. SIREN [26] addressed this by initializing the input frequencies uniformly in a range. While this ensures a bandlimit at the start of training, as it progresses, the layer composition introduces higher frequencies, leading to noisy reconstructions. We avoid this by providing a novel initialization for the first layer coupled with a bounding scheme which gives controls for limiting the MLP spectrum.
Other works [10, 5] employed MFNs [7] to control the bandlimit by applying a filter on the spectrum. However, this strategy usually introduces reconstruction artifacts since MFNs hard-truncate the spectrum. BANF [24] employed a grid-based MLP with a spatial filter that leverages grid resolution to constrain the highest frequency in the network. In our experiments, we observe that this is prone to creating higher frequencies which propagate as artifacts. In contrast, TUNER, grounded in a theoretical result (Thrm 2), guarantees a bandlimited MLP, and serves as a soft filter, providing a representation without ringing artifacts.
3 Sinusoidal MLPs as Fourier series
This work addresses the problem of deriving an efficient training scheme with controlled bandlimit for sinusoidal MLPs. This section presents the mathematical definitions and novel formulas for approaching this task.
Sinusoidal MLPs demonstrated high representational capacity with only a few hidden layers [26, 35, 19]. To understand how layer composition encodes this capacity, we propose to thoroughly investigate the structure of a $3$-layer sinusoidal MLP $f\!:\!\mathbb{R}^{d}\!\to\!\mathbb{R}$. For simplicity, we assume the codomain to be $\mathbb{R}$, however the same analysis holds for dimension $>\!\!1$. More precisely, we consider $f(\textbf{x})\!=\!\textbf{C}\!\circ\!\textbf{S}\!\circ\!\textbf{D}(\textbf{x})\!+\!e,$ with $\textbf{D}(\textbf{x})\!=\!\sin(\omega\textbf{x}\!+\!\varphi)$ being the input layer that projects x into a list of harmonics (input neurons) ${D}_{i}(\textbf{x})\!=\!\sin\left(\omega_{i}\textbf{x}+\varphi_{i}\right)$ with frequencies $\omega\!=\!(\omega_{1},\ldots,\omega_{m})\!\in\!\mathbb{R}^{m\times d}$ and shifts $\varphi\!\in\!\mathbb{R}^{m}$. Layer D is then composed with $\textbf{S}(\textbf{x})\!=\!\sin(\textbf{W}\textbf{x}+\textbf{b})$, where $\textbf{W}\in\mathbb{R}^{n\times m}$ is the hidden matrix, and $\textbf{b}\!\in\!\mathbb{R}^{n}$ the bias. Finally, $\textbf{C}\cdot\textbf{x}+e$ is an affine transformation with $\textbf{C}\in\mathbb{R}^{n\times 1}$ and $e\in\mathbb{R}$.
We now present some properties of the sinusoidal layers, which play key roles in the representation, and give a reinterpretation of them in terms of the network parameters. First, note that the hidden neurons $\textbf{h}(\textbf{x}):=\textbf{S}\circ\textbf{D}(\textbf{x})$ are defined by composing the dictionary D with the hidden sinusoidal layer which results in a list with elements
$h_{i}(\textbf{x})=\sin\bigg{(}\sum_{j=1}^{m}{W}_{ij}\sin\left(\omega_{j}\textbf{x}+\varphi_{j}\right)+b_{i}\bigg{)}.$ (1)
The following is a key result of this work, which states that we can linearize a hidden neuron (1) as a sum of sines with frequencies and amplitudes determined by $\omega$ and W.
Theorem 1.
Each hidden neuron $h_{i}$ of a 3-layer sinusoidal MLP has an amplitude-phase expansion of the form
$h_{i}(\textbf{x})=\sum_{\textbf{k}\in\mathbb{Z}^{m}}\alpha_{\textbf{k}}\,\sin\big{(}\beta_{\textbf{k}}\,{\textbf{x}}+\lambda_{\textbf{k}}\big{)},$ (2)
where $\beta_{\textbf{k}}\!=\!\left\langle\textbf{k},\omega\right\rangle$, $\lambda_{\textbf{k}}\!=\!\left\langle\textbf{k},\varphi\right\rangle\!+\!b_{i}$, and $\alpha_{\textbf{k}}\!=\!\prod_{j}\!J_{k_{j}}(W_{ij})$ is the product of the Bessel functions of the first kind.
The Bessel functions $J_{k}$ appear in the Fourier series of $\sin\big{(}a\sin(x)\big{)}$ [2, Page 361] and Thrm 1 generalizes this result. See the proof and generalization in the Supp. Mat. Yüce et al. [34] present a similar expansion for neurons activated by polynomials, however, they do not derive an amplitude formula. Their expansion for sinusoidal neurons is restricted to a 3-layer MLP without bias, a critical parameter for Fourier representation. Ours incorporates bias and generalizes to deep MLPs.
We now list some consequences of (2). First, the layer composition introduces a vast number of frequencies $\beta_{\textbf{k}}$ depending only on $\omega$, and shifts given by the input shift $\varphi$ and the bias b. More precisely, truncating the expansion by summing over $\left|\textbf{k}\right|_{\infty}\!\!\leq\!\!B\!\in\!\mathbb{N}$111$\left|\textbf{k}\right|_{\infty}$ denotes $\max\{\left|k_{1}\right|,\ldots,\left|k_{m}\right|\}$. implies that each hidden neuron $h_{i}$ could learn $\frac{(2B+1)^{m}-1}{2}$ non-null frequencies. This frequency generation gives a novel explanation of why composing sinusoidal layers greatly increases the network capacity. We will explore (2) to define a robust initialization for the INR’s input neurons in Sec 4.1.
Secondly, note that the weights W fully determine the amplitude $\alpha_{\textbf{k}}$ of each harmonic $\sin\!\big{(}\beta_{\textbf{k}}\textbf{x}\!+\!\lambda_{\textbf{k}}\big{)}$. Thus, to derive our bounding scheme we need to focus only on W. Finally, we can derive a sine-cosine form of (2),
$h_{i}(\textbf{x})\!=\!\!\!\sum_{\textbf{k}\in\mathbb{Z}^{m}}\!\!{A}_{\textbf{k}}\sin(\beta_{\textbf{k}}\textbf{x})\!+\!{B}_{\textbf{k}}\cos(\beta_{\textbf{k}}\textbf{x}),$ (3)
with ${A}_{\textbf{k}}\!=\!\alpha_{\textbf{k}}\cos(\lambda_{\textbf{k}})$ and ${B}_{\textbf{k}}\!=\!\alpha_{\textbf{k}}\sin(\lambda_{\textbf{k}})$. Note that the generated frequencies are independent of the index $i$, thus the hidden neurons share the same harmonics: $\sin\big{(}\beta_{\textbf{k}}\textbf{x}\big{)}$, $\cos\big{(}\beta_{\textbf{k}}\textbf{x}\big{)}$. Since different combinations of the input frequencies may correspond to a single frequency, (3) isn’t (yet) the Fourier transform of the network. In other words, we could have two integer vectors $\textbf{k},\textbf{l}\in\mathbb{Z}^{m}$ such that $\beta_{\textbf{k}}=\beta_{\textbf{l}}$. In the Supp. Mat., we show how to aggregate those frequencies to obtain the final Fourier transform.
To control the MLP’s bandlimit, we need a method to bound the amplitudes of $\beta_{\textbf{k}}$. Precisely, for a given number $B>0$ we would need $\left|\alpha_{\textbf{k}}\right|$ to be small for $\left|\beta_{\textbf{k}}\right|\geq B$. For this, we use the formula of $\alpha_{\textbf{k}}$ to determine a bounding:
Theorem 2.
The magnitude of the amplitudes $\alpha_{\textbf{k}}$ in the expansion (2) is bounded by $\prod_{j=1}^{m}\Big{(}\tfrac{|W_{ij}|}{2}\Big{)}^{|k_{j}|}\frac{1}{|k_{j}|!}.$
Sec 4.2 presents a bandlimit control mechanism during training based on Thrm 2.
Figure 2: Overview of TUNER. To train a sinusoidal MLP (gray model, top-left), we employ two techniques derived from Thrms 1 and 2. First, we initialize the input frequencies $\omega$ (green, bottom-left) with a dense distribution of low frequencies (red square) and a sparse distribution of higher frequencies (green grid). This initialization gives flexibility to learn the remaining signal frequencies which are simply integer combinations of $\omega$ (the yellow nodes on the right), a consequence of the amplitude-phase expansion given by Theorem 1. Note that this initialization resembles a frequency sampling since the training generates those new frequencies around $\omega$. Second, we bound the coefficients of the hidden layer weights (blue nodes) to ensure that the MLP remains within a specified bandlimit. This approach is effective because the amplitude-phase expansion (shown on the right) of each hidden neuron (purple nodes) indicates that the amplitudes of the generated frequencies have an upper-bound depending only on the hidden weights (blue, bottom-right).
4 Initialization and frequency bounding
We present TUNER, an initialization and frequency control scheme for sinusoidal INRs. Our initialization considers integer input frequencies, based on a Fourier series interpretation, and uses Thrm 1 to study the spectrum of layer composition. Then, we use Thrm 2 to control the INR’s bandlimit during training. Fig 2 provides an overview of TUNER. Throughout this section we present toy experiments to motivate the method, using $f\!:\!\mathbb{R}^{2}\!\to\!\mathbb{R}^{3}$ in an image reconstruction setup; later, comprehensive experiments will be given in Sec 5. We report the MLP size using the parameters $m$, $n$ as they completely determine the model. We use the Kodak dataset’s images [1] and train the INR using Adam [9].
4.1 Initialization as spectral sampling
A key challenge in initializing a sinusoidal MLP lies in defining its input frequencies $\omega$. Recall that the MLP generates frequencies as integer combinations of $\omega$, i.e., $\beta_{\textbf{k}}\!=\!\sum_{i}k_{i}\omega_{i}$, with the remaining weights representing their amplitudes $\alpha_{\textbf{k}}$. This setup presents two main challenges. First, the initialization of $\omega$ must enable $\beta_{\textbf{k}}$ to cover the signal’s spectrum; otherwise, the optimization cannot learn missing frequencies. Fig 3 (left) shows an example where using only odd frequencies leads to poor reconstruction. The second problem is the lack of flexibility for learning new frequencies. For example, randomly initializing $\omega$ may not be enough since it may produce many high frequencies in order to overfit the signal, see Fig 4 (top).
Figure 3: Choosing $\omega$ as the cartesian product of the odd frequencies without (left) or with (right) the frequencies $(1,0),(0,1)$. Note that adding them prevents sub-periods (see Supp. Mat.). We trained for $3000$ epochs on a $256^{2}$ image with network parameters $m=72$, $n=512$, and $\mathscr{b}=30$.
Figure 4: Uniform init. of $\omega$ (top) and ours (bottom). Grayscale bands show INR’s gradient. Ours offers better signal/gradient reconstruction w/o gradient supervision. The MLPs $(m\!=\!128,n\!=\!256)$ with bandlimit $\mathscr{b}\!=\!87$ were trained for $3000$ epochs.
To address these challenges, we introduce a novel initialization. First, we restrict our MLP to periodic functions, ensuring the training domain is in a full period. Such functions can be represented by sums of sines with integer frequencies, defining an orthogonal basis (Fourier series); This aligns well with our analysis (Thrm 1). The main challenge is initializing $\omega$ such that the generated frequencies $\beta_{\textbf{k}}$ cover the full spectrum within a bandlimit $B$ (Fourier basis).
More precisely, we initialize the weights $\omega$ with integer frequencies, that is $\omega_{j}\!\!\in\!\!\frac{2\pi}{p}\mathbb{Z}^{d}$, and freeze them during training, with $p\!>\!0$ being the period of the INR. This guarantees that the input neurons are $p$-periodic and Thrm 1 ensures that such periodicity is preserved over layer composition. Additionally, (3) implies that we can rewrite the INR as:
$f(\textbf{x})\!=\!\!\sum_{\textbf{k}\in\mathbb{Z}^{m}}\!\!\left\langle\textbf{C},{A}_{\textbf{k}}\right\rangle\sin\big{(}\beta_{\textbf{k}}\textbf{x}\big{)}\!+\!\left\langle\textbf{C},{B}_{\textbf{k}}\right\rangle\cos\big{(}\beta_{\textbf{k}}\textbf{x}\big{)}\!+\!e.$ (4)
Since the generated frequencies $\beta_{\textbf{k}}$ only depend on the frozen parameters $\omega$, the training is responsible for learning the amplitudes of the sine-cosine series in (4).
Let $B$ be the signal’s bandlimit. To ensure that the generated frequencies do not bypass $B$, we cannot sample $\omega$ directly on $[-B,B]^{d}$ since the MLP generates multiples of these frequencies. Therefore, we sample $\omega_{j}$ in $\frac{2\pi}{p}[-\mathscr{b},\mathscr{b}]^{d}$, with $\mathscr{b}\in(0,B)$ being a threshold for the input frequencies. Indeed, as the coefficients of hidden matrix W are limited by $2$ (see Sec 4.2), the magnitudes of ${A}_{\textbf{k}}$ and ${B}_{\textbf{k}}$ decrease as $\left|\textbf{k}\right|_{\infty}$ increases. This is a consequence of Thrm 2:
$|\alpha_{\textbf{k}}|\leq\prod_{j=1}^{m}\!\frac{\Big{(}\frac{|W_{ij}|}{2}\Big{)}}{|k_{j}|!}^{|k_{j}|}\!\!\!\!\!\leq\!\frac{1}{|k_{1}|!\ldots|k_{m}|!};\,\,i=1,\ldots,n.\vspace{-0.1cm}$ (5)
Thus, the generated frequencies $\beta_{\textbf{k}}$ with small k have more influence over the expansion (4), mimicking Fourier series.
For example, if input frequencies include 1 and 20, frequencies like 19 = 20 - 1 and 21 = 20 + 1 can be represented through their combinations. Thus, low frequencies give flexibility in expanding around the input ones which motivates initializing more input frequencies near the origin. We found $\mathscr{b}=\frac{B}{3}$ to work well experimentally. Fig 5 shows that it is necessary, as higher $\mathscr{b}$ makes it harder to learn the small frequencies, introducing noise.
Figure 5: Image reconstructions with bandlimit $B\!\!=\!\!128$, varying $\mathscr{b}\!=\!4,43,128$ with a network ($m\!=\!80,n\!=\!1000$) trained over 3000 epochs. The gradient magnitude is shown on the left of each image. Note that smaller $\mathscr{b}$ (middle-left) yield better reconstructions, while higher values of $\mathscr{b}$ introduce noisy gradients (right).
To initialize $\omega\!\in\!\mathbb{R}^{m\times d}$, we need to choose $m$ frequencies in $[-\mathscr{b},\mathscr{b}]^{d}$. Taking into account the importance of the lower frequencies discussed above, we split it into regions $\textbf{L}\!=\![-\mathscr{l},\mathscr{l}]^{d}$ of low and $\textbf{H}=[-\mathscr{b},\mathscr{b}]^{d}\setminus\textbf{L}$ of high frequencies, $\mathscr{l}\in(0,\mathscr{b})$. As $\beta_{\textbf{k}}$ only has a significant amplitude when $|\textbf{k}|_{\infty}$ is small (5), initializing with only low input frequencies may not allow the generation of high frequencies. Thus, we initialize some frequencies in H to cover the spectrum, and the rest in L to cover a neighborhood of the input frequencies. Specifically, we select $m_{l}<m$ frequencies uniformly from L. Then, we choose $m-m_{l}$ frequencies from H, spread in a grid-like fashion to cover $[-\mathscr{b},\mathscr{b}]^{d}$. In our experiments we considered $m_{l}=0.7\,m$. Finally, since $\textbf{D}(\textbf{x})\!=\!\sin(\omega\textbf{x}\!+\!\varphi)$, we initialize $\varphi$ uniformly in $[-\frac{\pi}{2},\frac{\pi}{2}]$.
When $\omega_{j}\!\in\!\textbf{L}$, small variations in $k_{j}$ results in a frequency near $\beta_{\textbf{k}}$. Indeed, Fig 6 shows how low frequencies (green square) introduce frequencies around the input frequencies (Fig 6a). Conversely, if $\omega_{j}\in\textbf{H}$ then $\beta_{\textbf{k}+\textbf{e}_{j}}$ is far away from $\beta_{\textbf{k}}$; we note this in Fig 6b-d, where $\omega_{j}$ is the red point and the arrows show the generated frequencies $k_{j}\omega_{j}$.
Figure 6: INR spectrum during training. The input frequencies $\omega$ in green. During training new frequencies appear in a neighborhood of $\omega$. The red dot indicates a high input frequency $\omega_{j}$ and the red arrows show the generated frequencies $k_{j}\omega_{j}$, $k_{j}=1,2,3$.
Finally, when sampling frequencies in L, we include $(1,0)$ and $(0,1)$ to guarantee that any frequency could be generated and prevents the appearance of sub-periods as shown at the beginning of this section in Fig 3 (left) (see also Supp. Mat for more details). Fig 3 (right) shows how this problem is solved using our initialization.
The proposed initialization overcomes the noisy reconstruction issues that arise with naive uniform sampling in $[-\mathscr{b},\mathscr{b}]^{d}$, as used in [26] (Fig 4 (top)). Fig 4 (bottom) demonstrates that our scheme significantly improves both signal and gradient reconstruction quality.
4.2 Bounding and hidden layer initialization
As shown in previous sections, sinusoidal INRs can represent detailed signals due to the wide number of frequencies generated by layer composition. However, they could manifest as noise in the reconstruction, see Fig 5 (right). Specifically, high input frequencies may generate even higher ones, bypassing the bandlimit $B$. To avoid this, we use Thrm 2 to introduce a bounding scheme to limit the amplitudes. Precisely, Thrm 2 implies that in each hidden neuron $h_{i}$, the amplitude of a generated frequency $\beta_{\textbf{k}}$ is limited by:
$|\alpha_{\textbf{k}}|\leq\left(\dfrac{\left|\textbf{W}_{i}\right|_{\infty}}{2}\right)^{\left|\textbf{k}\right|_{1}}\frac{1}{\prod_{j=1}^{m}|{k_{j}}|!}.$ (6)
[…]
Figure 7: Training a network $f$ with $m=n=256$ for $6000$ epochs with $\mathscr{b}\!\!=\!\!30$ and bounds $\mathscr{c}\!\!=\!\!0.1,0.2,0.5$. The bound over the hidden matrix W restricts the appearance of high frequencies.
In Sec 4.1 we split the input frequencies $\omega$ in two sets: the low ($\left|\omega_{j}\right|_{\infty}\!\leq\!\mathscr{l}$) and high ($\left|\omega_{j}\right|_{\infty}\!>\!\mathscr{l}$) frequencies. Additionally, we saw that the lower ones have a main role in the frequency generation, thus, we consider a higher threshold for them. For this, observe that each input frequency $\omega_{j}$ is associated with the $j$-column of W. Then, bounding the columns of W related to high frequencies, we restrict the amplitudes of $\beta_{\textbf{k}}$ that could exceed the Nyquist limit $B$. Conversely, since low frequencies are used to span around $\omega$, we consider a higher bound for their columns. Precisely, we define two parameters $\mathscr{c}_{\textbf{L}}$ and $\mathscr{c}_{\textbf{H}}$ to bound the weights $W_{ij}$ corresponding to $\omega_{j}$ in L and H, respectively,
$\left|\omega_{j}\right|_{\infty}\!\leq\!\mathscr{l}\,\,\longrightarrow\,\,W_{ij}=\texttt{clamp}\big{(}W_{ij},[-\mathscr{c}_{\textbf{L}},\mathscr{c}_{\textbf{L}}]\big{)},$
$\left|\omega_{j}\right|_{\infty}\!>\!\mathscr{l}\,\,\longrightarrow\,\,W_{ij}=\texttt{clamp}(W_{ij},[-\mathscr{c}_{\textbf{H}},\mathscr{c}_{\textbf{H}}]\big{)}.$
We use these bounds to initialize W such that each column is normal distributed with mean 00 and standard deviation $\tfrac{\mathscr{c}_{*}}{3}$, with $\mathscr{c}_{*}$ being the bound corresponding to that column. Note that SIREN [26] initializes $\textbf{W}\in\mathbb{R}^{n\times m}$ uniformly in $\small{\big{(}\!\!-\!{\sqrt{{6}/{m}}},{\sqrt{{6}/{m}}}\big{)}}$, and as $\small{\left|\textbf{W}\right|_{\infty}\!\leq\!\sqrt{{6}/{m}}<2}$ for $m>1$, it is bandlimited only at the beginning of training, since W is not controlled in any way during the optimization. Conversely, our initialization controls the bandlimit during all training and has faster convergence.
Additionally, we present a scheme to learn the bounding parameters during training. For this, we bound each $j$-column of the hidden matrix W using a $\tanh$ activation followed by multiplication with a trainable bounding parameter $\mathscr{c}_{j}$. Clearly, the resulting column is bounded by $\mathscr{c}_{j}$. In other words, we replace the hidden layer $\sin(\textbf{W}\textbf{x}+\textbf{b})$ by
$\small\sin\big{(}\tanh(\textbf{W})\mathscr{C}\textbf{x}+\textbf{b}\big{)},$ (7)
where $\mathscr{C}$ is a diagonal matrix with each $jj$-entry being $\mathscr{c}_{j}$. To prevent the bounds from growing too large, we use Thrm 2 to define the regularization term as $\mathcal{L}_{\textrm{reg}}=\sum|\mathscr{c}_{j}|$. Otherwise, the hidden weights may grow unbounded, leading to the INR’s overfit as in the usual training.
5 Experiments
This section presents experiments regarding the initialization and bounding of frequencies of an INR $f$ considering the images from Kodak dataset [1] resized to $[-1,1]^{2}$. We describe the architecture of $f$ with hidden matrix $\textbf{W}\!\in\!\mathbb{R}^{n\times m}$ with the parameters $m,n$. We initialize $f$ following the scheme in Sec 4.1 and optimize it using Adam [9] with learning rate $10^{-4}$. For the quantitative evaluations, we report the mean and standard deviation of the Peak Signal-to-Noise Ratio (PSNR) evaluated on the test set ($10\%$ random pixels). To show that TUNER adds high order regularity to the reconstruction, we compare the analytical gradient $\nabla f$ to the Sobel filter of the ground truth. All experiments were run in a 12 GB NVIDIA GPU (TITAN X Pascal).
5.1 Initialization
First, we compare our initialization of $\omega$ ($36.2$dB) against the uniform distribution ($32.2$dB). We obtain a $4$dB improvement of PNSR on average (with a std of $1.6$dB). We also present an ablation over the parameter $\mathscr{l}$ which splits the square $[-\mathscr{b},\mathscr{b}]^{2}$ in regions of low (L) and high (H) frequencies. Here we use $\mathscr{b}=85$ and $\mathscr{l}\!=\!21$, $42$, $64$, and sample $\omega$ with $30\%$, $50\%$, and $70\%$ of coordinates in H. Table 1 shows the performance of the INR.
[…]
$\mathscr{l}$ Signal Grad Signal Grad Signal Grad
$21$ 35.2 26.9 35.2 26.6 35.0 26.0
$42$ 35.1 26.3 34.8 24.6 33.8 22.2
$64$ 34.2 25.2 33.3 21.6 30.8 15.0
Table 1: Training an INR with $m=n=416$ parameters, for $400$ epochs. Results indicate that it is better to sample more $\omega_{j}\in\textbf{L}$ and choose a smaller $\mathscr{l}$.
This experiment considered all the images in the dataset. Observe that all values of the row $\mathscr{l}\!=\!64$ have lower PSNR, showing that an INR with very high frequencies performs worse even for small percentages ($30\%$). Similarly, the performance worsens when we increase the number of high frequencies (last column).
We now test the influence of high frequencies by training an INR of size $m\!=\!360$, $n\!=\!512$ with $\mathscr{b}\!=\!256$ over $90\%$ of the ground truth pixels. We vary the parameter $\mathscr{l}\!=\!60,120,220$ that defines low frequencies. Increasing the number of higher frequencies, the INR suffers from noise and overfitting. This can be noted in the reconstruction on the unsupervised pixels and gradients in Fig 8. Finally, in Supp. Mat. we compare our initialization of W against SIREN and observe a faster convergence.
Figure 8: Reconstructions with big $\mathscr{l}$ present overfitting on the $10\%$ unsupervised points. Gradient on the upper right corner.
5.2 Bounding
In Sec 4.2 we presented a scheme to bound the INR spectrum. Here, we show additional experiments to demonstrate that this improves the signal and has a great benefit over the unsupervised gradient reconstruction. Then, we give a comparison between fixed and trained bounding schemes.
Fig 9 compares the reconstructions of an image of size $512^{2}$ wo/w our fixed bounding scheme. To demonstrate the impact of our scheme, we sample the input frequencies using $\mathscr{b}\!=\!256$ and $\mathscr{l}\!=\!100$. Without bounding (left), we effectively reconstruct the signal at the supervised pixels (30.8 dB), but with a noisy gradient (15.8 dB). That is, we are overfitting the signal. Conversely, bounding the hidden weights significantly improves both the signal (35.1 dB) and its gradient (27.8 dB). Our bounding scheme acts like a filter in both signal and gradient reconstruction.
Figure 9: Comparison of training wo/w fixed bounds. Note that the use of bounds preserves high order information of the signal. Left side of the figures show the network gradients.
We compare quantitatively our bounding scheme over the dataset, varying the parameters $\mathscr{b}$ and $\mathscr{c}_{\textbf{H}}$. Table 2 shows an improvement in the signal for all cases. Furthermore, the unsupervised gradient of the image is better learned when using fixed bounds, showing an average increase of 9.6dB against the standard training. Also, the last two lines of the table show the impact of the parameters $\mathscr{c}_{\textbf{L}}$, $\mathscr{c}_{\textbf{H}}$ over the signal and gradient reconstructions. We consider training with $\mathscr{c}_{\textbf{L}}>\mathscr{c}_{\textbf{H}}$ since we are prioritizing low input frequencies in the spectrum generation. Otherwise, this may generate high frequencies, possibly bypassing the Nyquist limit and resulting in a worse gradient reconstruction.
signal grad
$\mathscr{b}$ $\mathscr{c}_{\textbf{L}}$ $\mathscr{c}_{\textbf{H}}$ wo bound w bound wo bound w bound
256 1.5 0.05 32.6 34.8 14.5 25.8
190 1.5 0.05 34.0 35.9 18.5 26.1
190 1.5 0.2 34.0 34.4 18.5 21.0
Table 2: Training with bounding improves the signal and grad. reconstruction. The choice of the fixed bound $\mathscr{c}_{\textbf{H}}$ greatly impacts the grad. fidelity. The std of the difference between each image signal (grad) quality wo/w bounding is $1.4$dB (1.6dB) on average. Evaluations performed on the test set ($10\%$ of pixels).
Next, we show how we can learn each bound $\mathscr{c}_{j}$. As described in Sec 4.2, we can incorporate a learnable $\mathscr{c}_{j}$ for each $j$-column of the hidden matrix through a modified layer (7). Fig 10 shows a comparison between our fixed bound scheme using initial un-tuned $\mathscr{c}_{\textbf{L}}\!\!=\!\!1.0$ and $\mathscr{c}_{\textbf{H}}\!\!=\!\!0.6$ (left) and the learned bound scheme (middle).
Figure 10: Comparison of signal and gradient (grayscale) when training with fixed/learned bounds. The learned bounds adjust the INR spectrum resulting in better reconstructions. Each blue point represents a pair $\left(|\omega_{j}|_{\infty},\mathscr{c}_{j}\right)$, where $c_{j}$ is the trained bound of $\omega_{j}$.
As expected, training lowers the bounds for high frequencies (H) while maintaining or increasing the bounds of low frequencies (L). Thus, we achieve better reconstruction in both the signal and gradient. In Fig 10 (right), we note that low frequencies (left of the green line) have significant bounds, showing their importance during training. These schemes allow us to initialize with higher frequencies than $\nicefrac{{B}}{{3}}$, accelerating convergence. Finally, we performed an ablation study on the regularization term $\mathcal{L}_{reg}$ to show its importance. Training w/wo regularization yielded $28.5/27.4$dB for the RGB and $27.1/24.0$dB for the gradient.
5.3 Comparisons
We first compare our method (TUNER) with SIREN, which requires choosing a threshold parameter $\mathscr{b}$, [^2] Denoted as $\omega_{0}$ in [26]. a critical task since higher $\mathscr{b}$ implies noisy reconstructions. TUNER also uses $\mathscr{b}$ to control the frequency generation, however, it provides more stable training with better reconstruction. To demonstrate this, we perform a comparison by varying $\mathscr{b}$ and the number of input ($m$) and hidden ($n$) neurons. First, consider the case where we have no information about the signal bandlimit besides the Nyquist limit $B$. Fig 11 presents a comparison between SIREN (top) and TUNER (bottom) with $\mathscr{b}=B$. Note that SIREN produces noisy reconstructions, while ours learns faster and does not suffer from noise. This is due to our bounding scheme, which limits the generation of high frequencies and provides stable training (Sec 4.2). In contrast, SIREN initializes its parameters such that the spectrum is initially bounded but does not maintain this guarantee during training.
Figure 11: Comparing SIREN (top) and TUNER (bottom) when we only know the Nyquist limit (i.e, $\mathscr{b}\!=\!B$). Since $\mathscr{b}$ is big, SIREN introduces high frequencies at the beginning of training, resulting in very noisy reconstructions. Conversely, ours starts training low frequencies, converging faster and being robust to $\mathscr{b}$.
Next, we consider a tuned $\mathscr{b}$ for SIREN to show that even under this condition, our method provides a better training scheme for sinusoidal INRs. Fig 12 presents the comparison varying the number of input ($m$) and hidden ($n$) neurons. We consider the cases $m<n$ (top) and $m>n$ (bottom), and TUNER performed better both in the signal/gradient reconstructions. Also, we note more robust training using TUNER, as shown in Fig 12(right).
The previous experiment shows that initializing SIREN with a high $\mathscr{b}$ results in overfitting, while TUNER provides better reconstruction. On the other hand, with the same architecture and a tuned $\mathscr{b}$, our method also effectively learned the lower frequencies first and performed better (Fig 12).
Figure 12: Comparison with SIREN when $\mathscr{b}=40$. TUNER improves training convergence in both cases ($m<n$, $m>n$) and stability when $m>n$. INR gradient shown in grayscale.
Next, we numerically evaluate these experiments, providing a quantitative comparison with SIREN and BACON. We compare different architectures ($m$, $n$) and spectrum bandlimits $\mathscr{b}$, evaluating the results in terms of signal and gradient PSNR over the supervised (Train) and unsupervised (Test) pixels. Table 3 presents these evaluations.
Train Test
Signal Gradient Signal Gradient
$\mathscr{b}$ $(m,n)$ SIREN BAC ours SIREN BAC ours SIREN BAC ours SIREN BAC ours
85 1028,162 32.9 - 32.9 26.5 - 28.8 31.9 - 32.2 25.5 - 28.2
416,416 34.0 26.4 34.3 27.7 22.4 28.9 32.9 25.4 33.4 27.1 22.3 28.4
80,2024 32.2 - 33.1 26.6 - 28.5 31.3 - 32.3 26.0 - 28.0
171 1028,162 31.3 - 33.9 22.0 - 27.2 22.9 - 32.9 20.7 - 26.8
416,416 29.2 29.3 34.4 17.4 26.7 27.0 26.7 26.2 33.1 16.2 26.3 26.4
80,2024 26.6 - 30.8 14.2 - 22.6 23.8 - 29.5 13.8 - 21.3
256 1028,162 26.8 - 33.3 12.8 - 24.7 24.1 - 31.9 12.7 - 23.9
416,416 24.7 27.7 41.2 12.9 24.2 23.0 22.3 22.7 33.2 13.2 22.8 21.2
80,2024 25.1 - 30.0 13.0 - 15.2 24.1 - 27.1 13.1 - 14.6
Table 3: Comparison between SIREN, BACON (BAC), and our method. We use several architectures ($m$, $n$), bandlimits ($\mathscr{b}$), and evaluate over the train (90%) and test(10%) sets of signal/gradient.
As expected, SIREN performs well with $\mathscr{b}=\frac{B}{3}$. In this case, TUNER is perceptually similar to SIREN but improves gradient reconstruction and consistently outperforms BACON. When $\mathscr{b}=171$, TUNER offers superior quality in both signal and gradient, showing greater robustness to noise and superior performance even with $\mathscr{b}\geq\nicefrac{{B}}{{3}}$. Finally, with $\mathscr{b}=256$, our method surpasses SIREN in all cases. For BACON, our method achieves better signal reconstruction, with comparable gradients. However, the following experiment shows that BACON’s reconstruction has visible defects in this scenario. Note that we do not supervise the training with gradient information.
Bandlimit comparison:
Lastly, Fig 13 presents a qualitative comparison between BACON, BANF, and TUNER (Ours) with bandlimits $\mathscr{b}=60,130$. We train all methods with 6767~{}6767K parameters. For BANF we use the linear filter and adjust grid resolution to match the bandlimits and the MLP size to approximate $67$K parameters. Particularly, the spectra were computed with the same process for all three methods. Note that BACON truncates the spectrum with a hard filter, resulting in ringing artifacts (low bandlimit) and noisy reconstruction (high bandlimit). BANF shows distortions around the eye area for both filters, and the spectra show no clear bandlimit. Conversely, TUNER operates as a soft filter providing better reconstructions for all cases.
Figure 13: Comparison between BACON, BANF, and TUNER, trained with bandlimits $\mathscr{b}=60,130$. Note that BACON uses a box filter, generating ringing artifacts (left). BANF (middle) has artifacts near edges and ambiguous bandlimit. Then, TUNER (right) resembles a soft filter, improving quality.
6 Conclusions and limitations
We presented a study based on Fourier theory for sinusoidal MLPs, which resulted in novel, robust schemes for both initialization and spectral control of such networks. While our experiments focused on images to validate our theoretical claims, future work includes applying our schemes to other signal types. Additionally, we have focused our studies on 3-layer MLPs, leaving the investigation of deep networks for future research. We consider this work a first step toward unraveling the powerful expressiveness of sinusoidal MLPs.
Acknowledgments
We would like to thank CAPES, FAPERJ, and Google for partially funding this work.